Draw gravitational field lines to represent Jupiter’s gravitational field.

(b)Jupiter has a radius of 7.14 x 107m and the gravitational field strength at its surface is 24.9N/kg.

(i)Show that the mass of Jupiter is about 2 x 1027 kg.

(ii) Calculate the average density of Jupiter. 

1
Expert's answer
2021-02-26T10:23:00-0500

Answer

Draw gravitational field lines to represent Jupiter’s gravitational field as below



Now

b(i)

Mass can written as

"M=\frac{gr^2}{G}\\=\frac{24.9\times(7.14\times10^7)^2}{6.67\times10^{-11}}\\=2\times10^{27}Kg"

b(ii)

Density is written as

"\rho=\frac{3g}{4\pi RG}\\=\frac{3\times24.9}{4\times3.14\times 7.14\times 10^7\times 6.67\times10^{-11}}\\=1300kg/m^3"